OpenStudy (anonymous):

Having a hard time solving radical equations with two radicals. sqrt(4y+5)-sqrt(y-1)=3

OpenStudy (anonymous):

(4y+5)-(y-1)=9 3y=3 y=1

OpenStudy (anonymous):

I added sqrt(y-1) on both sides to get sqrt(4y+5)=3+sqrt(y-1) then squared both sides to get rid of the square roots to get 4y+5=(3+sqrty-1)squared but get lost from that point on.

OpenStudy (gw2011):

If you square both sides of the equation, you get: (4y+5)-(y-1)=9 4y+5-y+1=9 3y+6=9 3y=3 y=1

OpenStudy (anonymous):

What's all this??? Everybody forgot how to square things? mathisnofun started off right, continuing from there: 4y +5 = 9 + y -1 + 6 sqrt(y-1) (y'all call this FOIL, right?) --> y-1 = 2sqrt (y-1) and square again --> y^2 -6y + 5 = 0 --> (y-1)(y-5) --> y = 1 or 5.